OpenStudy (anonymous):

64s^2+144s=-81 can anyone help on this one. Solving for s

OpenStudy (anonymous):

64s^2+144s+81=0 (8s+9)^2=0 8s+9=0 8s=-9 s=-9/8

OpenStudy (anonymous):

hmmm?

OpenStudy (anonymous):

@marsss Thank you for the setup

OpenStudy (anonymous):

I think this problem is quadtradic

OpenStudy (anonymous):

i believe it is too

OpenStudy (anonymous):

@godorovg you're a genius :)

OpenStudy (anonymous):

@treydwg you're too :)

OpenStudy (anonymous):

thanks lol

OpenStudy (anonymous):

so the solution set would be

OpenStudy (anonymous):

I working on it

OpenStudy (anonymous):

in this problem the first clue is being = 0 usally says Quad needs to be used.

OpenStudy (anonymous):

What is the easier way of working out qUADRATIC without using the quad formula?

OpenStudy (anonymous):

64^2 is a peferct square so that always tells that this will invole in a square root to slove

OpenStudy (anonymous):

you could use a^2+B^2=c^2 as well

OpenStudy (anonymous):

So I got -1.125, -1.125. Can I write -1.125^2. Or is this considered incorrect?

OpenStudy (anonymous):

yes you can do that with 1.125^2 if you used the Quad form you would plus ninus which is what the problem is asking for..

OpenStudy (anonymous):

Ok i just wrote it twice. lets see if it counts it correct?

OpenStudy (anonymous):

treydwg first thing to know in this problem 64s^2+144s+81=0 all of the numbers are perfect squares. what that means taking the square root will be a number for example you take square root of 81 which is, we can check this by 9^2 and yes it equals 9.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about 2v^2+7v+5 and I need tofactor by grouping?

OpenStudy (anonymous):

I am at (2v^2+5v)(2v+5) and stuck

OpenStudy (anonymous):

@godorovg can you assist?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where do I go from here?

OpenStudy (anonymous):

in this problem we need to use F,.O.I .L

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

so now take the 2v^2 and times to 2v what is the answer

OpenStudy (anonymous):

4v^2

OpenStudy (anonymous):

close try again

OpenStudy (anonymous):

4v is right your power is wrong

OpenStudy (anonymous):

where are you getting the 2v^2*2v?

OpenStudy (anonymous):

because you typed 4v^2 while ago

OpenStudy (anonymous):

recall the problem here (2v^2+5v)(2v+5)

OpenStudy (anonymous):

yes that what I have

OpenStudy (anonymous):

2v(v+5)

OpenStudy (anonymous):

the (2v+5) is where im stuck

OpenStudy (anonymous):

what can I factor out of that?

OpenStudy (anonymous):

now 4v is right because 2v times 2v =4v^2 but that is not what is happening here. You have a power being times by V what is the power of the V? that power of the V is ^1 and when times 2v^2 time 2v = 4v^3

OpenStudy (anonymous):

so what happened here is we times the 2's to get 4 and v^2 + v^1 =V^3 see that

OpenStudy (anonymous):

let me look at that hold on

OpenStudy (anonymous):

ok I see it

OpenStudy (anonymous):

now how do I fix the last part?

OpenStudy (anonymous):

if you have a question plz ask me..

OpenStudy (anonymous):

factoring will not be in use unti we finish F.O.I.L

OpenStudy (anonymous):

I have (2v+5) how can I faactor this

OpenStudy (anonymous):

no you can not factor that..

OpenStudy (anonymous):

ok I see now so this is what I have after foil 4v^3+10v^2+35v

OpenStudy (anonymous):

now we can factor

OpenStudy (anonymous):

oooooooooh

OpenStudy (anonymous):

by grouping?

OpenStudy (anonymous):

no grouping there are no a like terms here

OpenStudy (anonymous):

I know but i was going to use 35v and try to make another number to group

OpenStudy (anonymous):

can not do that srry

OpenStudy (anonymous):

ok so wehre from here

OpenStudy (anonymous):

no GCF in common

OpenStudy (anonymous):

okay let us look at this 4v^3+10v^2+35v what is the same with all the numbers?

OpenStudy (anonymous):

v

OpenStudy (anonymous):

yes V so how many V's can be factored?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

v(2v^2+10v+35)

OpenStudy (anonymous):

now look at the numbers what can each number be divided by?

OpenStudy (anonymous):

v(4v^2+10v+35)

OpenStudy (anonymous):

that is the answer it can not be factored any further

OpenStudy (anonymous):

ahhhhhhh! Thats my problem i want to always factor further :/

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

treydwg here is a help for you, if you have done all the factoring right than find a number that divides in to the first and see if it works for example 2 will go ino 4 =2 2 will also go into 10 five times but here is a question for you will 2 divide even into 35?

OpenStudy (anonymous):

if the answer is no than you are finished..

OpenStudy (anonymous):

no it wont and that was my problem

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome anytime

OpenStudy (anonymous):

before I leave any questions plz ask me

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

working on another

OpenStudy (anonymous):

If you need help I am here

OpenStudy (anonymous):

5x-3=14x+6 I got x=-1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

5x-3=14x+6 get a like terms together so we x terms let us put the tgether like this 14x-5x=6-3 recall when a + or - crosses the equals sign it changes

OpenStudy (anonymous):

x=3/9??

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

ok i did that one a little wrong thank you

OpenStudy (anonymous):

now this one is throwing me off cuz the a^2 I have a^2=10-3a I got a^2=5/2

OpenStudy (anonymous):

no now think about this one? come you do it!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first off you can not a^2 = 10 -3a and come out with a^2=5/2 no..

OpenStudy (anonymous):

hint what happens to the problem when all nunbers are moved like this a^2+3a-10= what is this problem?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

what did I say when a problem is set to zero is?

OpenStudy (anonymous):

nothing is on one side

OpenStudy (anonymous):

or they are all on one side rather

OpenStudy (anonymous):

have to factor?

OpenStudy (anonymous):

recall the problem a^2 = 10 -3a now look at the problem a^2+3a-10=

OpenStudy (anonymous):

0

OpenStudy (anonymous):

we have to find something in common

OpenStudy (anonymous):

so what type of problem is this? not factoring

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i need to find out what two numbers product is 10 and sum is 3 right?